OpenStudy (anonymous):

find an equation of (-1,-2) perpendicular to 2x+5y+8=0

OpenStudy (texaschic101):

first we are going to put the equation in y = mx + b form... 2x + 5y + 8 = 0 2x + 5y = -8 5y = -2x - 8 y = -2/5x - 8/5 what we just did was find the slope of the equation. With the formula, y = mx + b, the m position is the slope, therefore the slope of this equation is -2/5. We are looking for a perpendicular line, so the slope we need is the negative reciprocal of the slope of the equation. All that means is " flip " the fraction and change the sign, and you will have the negative reciprocal. So the negative reciprocal of - 2/5 is 5/2. That is the slope we need. Now, since we have the slope and a set of points, we will use this formula : y - y1 = m(x - x1) slope(m) = 5/2 (-1,-2) x1 = -1 and y1 = -2 now lets sub our info into the formula : y - y1 = m(x - x1) y - (-2) = 5/2(x - (-1) y + 2 = 5/2(x + 1) y + 2 = 5/2x + 5/2 y = 5/2x + 5/2 - 2 y = 5/2x + 5/2 - 4/2 y = 5/2x + 1/2 <-- your equation